Giant pulses from energetic pulsars have been observed with S5GHz = 10 kJy, but with durations of only 10 ns. The region from where a pulse of radiation originates must be no larger than the distance that light can travel during the duration of a pulse. If the pulsars are at a distance of 1 kpc, estimate the brightness temperature of these sources.


1
Expert's answer
2022-03-30T03:31:34-0400

"T=\frac{c^2B}{2k\nu^2}=4.5\cdot 10^{35}~K."